OpenStudy (anonymous):

plase help! (x+3)^2-(x+3)-56 i have to factor completely

hero (hero):

x^2 + 6x+9 -x - 3-56 x^2 +5x +9 - 59 x^2 +5x - 50 x^2 + 10x - 5x - 50 x(x+10) - 5(x+10) (x+10)(x-5)

OpenStudy (anonymous):

first factor as \[((x+3)+7)(x+3)-8)\] then combine like terms in the parentheses to get \[(x+10)(x-5)\]

OpenStudy (anonymous):

-56=-8*7 and jus write x+3 where you would usually just write x, or make substitution so its more pleasing to the eye let a=x+3, then a^2-a-56=(a-8)(a+7) don't forget to sub back into x, if you do it this way

hero (hero):

Satellite, you're weird